Recent Log Entries

XSEP2T tripped off

Log Number: 4032624Logbook(s): Entry Makers: mockBody: 

Power supply screen did not show any faults.
Reset, turned on at setpoint2, ramped to setpoint 1, cycled by itself, then cycled with rest of XSEPs without any issue.

Box Supplies Were Cycled

Log Number: 4032623Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
XSEP2T Ramp Rate = 5.753 A/s
XSEP4T Ramp Rate = 5.314 A/s
XSEP6T Ramp Rate = 4.959 A/s
XSEP8T Ramp Rate = 17.478 A/s
XSEPAT Ramp Rate = 16.459 A/s

2L23-4,6 SEL quenches

Log Number: 4032592Logbook(s): Entry Makers: mockBody: 

Zone tripped and started auto recovering, the GLOS in cavities 4 and 6 were set too high causing the cavities to trip in SEL. This caused neighboring cavities to trip which were otherwise stable. After lowering the GLOS back down to 1.0 in cavities 4 and 6, was able to get cavities back into SEL and tune back in. Once all cavities looked good in SEL ramped them up in GDR.

Machine orbit

Log Number: 4032538Logbook(s): Entry Makers: mockBody: 

Hall B: 4 nA 5th pass
Hall C: 40 uA 3rd pass

COMMENT ON Add External DCCT to TAG5C

Log Number: 4032530Logbook(s): Body: 

Final cable swapped in, all hardware is in place. Waiting for work to finish in tagger before we ramp supply back up and verify "functionality". This functionality does not affect power supply, just the new external DCCT readback. Aiming to do this Wednesday but no urgency, could take place when supply comes back up as hall is preparing to take beam.

Operator SIGN IN

Log Number: 4032514Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Mock
DATE:             08/22/2022
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

Autolog save

Log Number: 4032510Logbook(s): Tags: Body: saver ran from opsl632

Saved: chumphry 08-22-22 16:58:13 {A:2 B:5 C:3 E:1047 A=BS B=CW C=CW D=BS 0R08=40.38 No dumps in.

Hall C -> Loop 2 set for 40 uA

Log Number: 4032502Logbook(s): Entry Makers: mockBody: 

Ion chambers set empircally due to Hall C configuration has the spectrometers at a small angle. Requested 40 uA due to rate limits on detectors, might change from this current.

Injector work summary

Log Number: 4032501Logbook(s): Entry Makers: ywanghoflerBody: 

Problem:
Hall B current spikes over 25nA when discharging in 0L02 region hapens and the requested current is 4nA.

Action:
1. Lowering prebuncher amplitude from 3.2 to 3.0
2. Adjusted Hall B slit from 17.0 to 16.5
3. Some minor steering in the chopper region

Will monitor to see if this mitigates the hall current spikes.

Autolog save

Log Number: 4032497Logbook(s): Tags: Body: saver ran from opsl06

Saved: ywang 08-22-22 16:28:13 {A:2 B:5 C:3 E:1047 A=BS B=BS C=Tune D=BS 0R08=5.67 dump(s) in.

Pages